OpenStudy (anonymous):

Simplify the radical expression √15y*3√81y A.2√135y^2 B.2y√135 C.27√15y^2 D.27y√15

OpenStudy (jdoe0001):

\(\large \sqrt{15y^3}\cdot 3\sqrt{81y}?\)

OpenStudy (anonymous):

\[\sqrt{15y}*3\sqrt{81y}\]

OpenStudy (jdoe0001):

\(\large \bf \sqrt{15y}\cdot 3\sqrt{81y}\implies 1\sqrt{15y}\cdot 3\sqrt{81y}\implies 3\cdot 1\sqrt{15y\cdot 81y}\)